OpenStudy (anonymous):

HELP How would you put the words "Inverse-variation power function" and "Inverse-square-variation power function" into actual functions???

OpenStudy (anonymous):

\[y = kx^a\] power function assuming inverse means what i think it means \[x = ky^a\] solve for y \[y = (\frac{x}{k})^{1/a}\]

OpenStudy (anonymous):

square power function would be x^2, i think? so a =2

OpenStudy (anonymous):

uh... idk...imma ask teacher l8r~~~ thanx!

hartnn (hartnn):

inverse variation means \[\huge y=\frac{k}{x}\] so, inverse power will be \[\huge y=\frac{k}{x^a}\] then for inverse square, just put a=2 \[\huge y=\frac{k}{x^2}\]

OpenStudy (anonymous):

a=2? why?

hartnn (hartnn):

square of x is x^2 so, inverse square variation should have exponent =2

OpenStudy (anonymous):

ohhhhh~ ok! Thank you!

hartnn (hartnn):

welcome ^_^